Portfolio returns. The Capital Asset Pricing Model is a financial model that assumes returns on a portfolio are normally distributed. Suppose a portfolio has an average annual return of 16.5% (i.e. an average gain of 16.5%) with a standard deviation of 32%. A return of 0% means the value of the portfolio doesn't change, a negative return means that the portfolio loses money, and a positive return means that the portfolio gains money. Round all answers to 4 decimal places.

a. What percent of years does this portfolio lose money, i.e. have a return less than 0%?

b. What is the cutoff for the highest 19% of annual returns with this portfolio?
Answer #1

a.

The percent of years does the portfolio lose money. That is, find the probability P(X<0)

Let X be the random variable defined by returns on a portfolio follows normal distribution with mean(μ) 16.5 % and standard deviation(σ) 32%.

The probability P(X<0) is,

P(X<0)=P(x-16.5/32<0-16.5/32)

=P(z<-0.52)

=0.3031 ============>> Use ''=NORMSDIST(-0.52)'' in Excl.

The percent of the year’s portfolio lose money, that is have a return less than 0% is 30.31%

b.

The cutoff for the highest 19% of annual returns with this portfolio is obtained below:

P(X>x)=0.19

P(X<x)=1-0.19

=0.81

From the “standard Normal table”, the area covered for value of 0.81 is obtained at . z=0.88

The cutoff for the highest 19% of annual returns with this portfolio is,

Z=(x-mean)/standard deviation.

0.88=x-16.2/32

x=16.5+0.88*32

=44.66

The cutoff for the highest 19% of annual returns with this portfolio is 44.66%

Hope this will be helpful. Thanks and God Bless You:)